# When Collaborative Maintenance Falls Short: The Persistence of Retracted Papers on Wikipedia

### v2

@@ 5.1.3 Automating Credibility Cues through Structured Metadata
While editors can manually apply retraction templates, there is no consistent or automated mechanism to propagate retraction warnings across articles that cite the same paper. Our study reveals that corrections are uneven and often delayed, sometimes by years, resulting in gaps in credibility signaling. To reduce this burden, citation templates could be expanded to automatically ingest and display structured metadata from Crossref, Retraction Watch, or similar sources. If a DOI is retracted, the template could render a default warning, either inline or as a footnote, linking to the retraction notice. This would both standardize correction practices and reduce the cognitive load for editors.
@@ 5.1.4 Surfacing External Signals for Epistemic Risk
Finally, we find that external signals of visibility and accessibility are linked to correction speed. Retracted papers mentioned more frequently on Twitter/X are corrected significantly faster, while paywalled papers are corrected more slowly. This suggests that social media attention functions as an external signal, surfacing now-invalidated citations, while a lack of open access may act as a barrier to verification and repair. Future tools might incorporate external alerting systems, integrating tweet activity and public discussions on platforms such as blogs or PubPeer regarding specific retractions into internal Wikipedia tools. This could help editors recognize emerging credibility issues even in articles that lack sustained editorial engagement.
@@ 5.2 Limitations and Future Work
Second, although our modeling identifies significant predictors of the speed with which retracted citations are corrected, it cannot fully account for the editorial processes and sociotechnical dynamics that shape citation repair on Wikipedia. Decisions to add, retain, or remove citations are often entangled with disputes over content reliability, page notability, or conflicting interpretations of sourcing policies. Our analysis does not surface these nuanced deliberations, which may vary across topics, editor groups, or article histories. Qualitative research—such as interviews with active editors, analysis of talk page discussions, or observational studies of how citation templates are used—could provide deeper insight into how retraction cues are interpreted and acted upon in practice.
Another limitation lies in the scope of our survival analysis model (specifically, the Weibull AFT model), which, while effective for identifying associations, does not capture the complex interactions between features. For instance, factors such as citation timing, editorial activity, and retraction reasons may interact in non-linear ways. Future work could employ more sophisticated machine-learning approaches or mixed-methods frameworks to uncover deeper patterns and relationships in the data.
@@ 6 Conclusion
This study examined the presence and persistence of retracted paper citations on Wikipedia, identifying key factors influencing the speed of corrective actions. Our findings reveal that while some citations are corrected promptly, many persist uncorrected for extended periods, with a median time to correction of over 3.68 years (1,344 days). The speed of correction is significantly accelerated by factors such as a salient retraction reason, high Twitter/X attention, open access publication, a greater number of unique Wikipedia editors, and whether the paper was initially noted as controversial. Conversely, a high academic citation count is associated with a slower correction time. These findings highlight a fundamental disconnect between the public availability of retraction data and the community’s collaborative workflows, which often fail to translate this information into the reader-facing edits required for a complete repair. Addressing this challenge requires moving beyond simple error flagging towards a tighter integration of information systems and human editorial processes. Design interventions that surface different types of signals to support editorial work—such as contextual signals embedded in editing interfaces, structured data signals within citation templates, and prioritized attention signals on community dashboards—could significantly improve responsiveness to invalidated science. By combining better information routing with sustained human oversight, Wikipedia can more effectively manage citation-level credibility failures and strengthen its role as a reliable public knowledge platform.
@@ Appendix A Survival Analysis Model Selection Details
Table A1: Test of the Proportional Hazards Assumption. The test is based on Schoenfeld residuals. A significant p-value (p < 0.05) indicates that a variable violates the assumption.
Variable Chi-Square ($\chi^{2}$) df p-value
Open Access 0.18 1 0.67
Academic Citation Count (at risk start) 0.02 1 0.88
Team Size 1.88 1 0.17
Physical Sciences (domain) 0.47 1 0.49
Life Sciences (domain) 4.46 1 0.03
Social Sciences (domain) 0.24 1 0.62
Tweet Mention Count 1.10 1 0.29
Time to Retraction (days) 0.73 1 0.39
Pre-Retraction Editorial Event 1.01 1 0.32
Retraction Reason (salient) 1.21 1 0.27
Unique Editors (at risk start) 2.37 1 0.12
Controversy Noted Initially 0.24 1 0.63
Table A2: Comparison of Accelerated Failure Time (AFT) Model Fit. The model with the lowest Akaike Information Criterion (AIC) was selected for the final analysis.
Model Distribution AIC Concordance Log-Likelihood Ratio Test ($\chi^{2}$)
Weibull 8209.512 0.671 117.551
Log-Logistic 8310.896 0.676 134.926
Log-Normal 8478.694 0.673 130.386
Appendix B Feature Correlation and Multicollinearity
Figure B1: Correlation matrix of independent variables used in the Weibull AFT model. Most features exhibit low to moderate correlation, with no evidence of severe multicollinearity.
Table B1: Variance inflation factors (VIFs) for all independent variables. All VIFs are well below the common threshold of 5, indicating no multicollinearity.
Feature VIF
Open Access 1.09
Academic Citation Count (at risk start) 1.93
Team Size 1.34
Physical Sciences (domain) 1.27
Life Sciences (domain) 1.47
Social Sciences (domain) 1.15
Time to Retraction (days) 1.81
Pre-Retraction Editorial Event 1.24
Retraction Reason (salient) 1.20
Tweet Mention Count 1.49
Unique Editors (at risk start) 1.19
Controversy Noted Initially 1.05

### v3

@@ 5.1.3 Automating In-Text Credibility Cues through Structured Metadata
While editors can manually apply retraction templates, there is no consistent or automated mechanism to propagate retraction warnings across articles that cite the same paper. Our study reveals that corrections are uneven and often delayed, sometimes by years, resulting in gaps in credibility signaling. To reduce this burden, citation templates could be expanded to automatically ingest and display structured metadata from Crossref, Retraction Watch, or similar sources. If a DOI is retracted, the template could render a default warning, either inline or as a footnote, linking to the retraction notice. This would both standardize correction practices and reduce editors’ cognitive load. One promising route is to leverage community-maintained structured sources within the Wikimedia ecosystem (e.g., Wikidata: WikiProject Retractions) as an intermediary layer for propagating retraction status across pages and interfaces.
@@ 5.1.4 Surfacing External Signals for Epistemic Risk
Finally, we find that external signals of visibility and accessibility are linked to correction speed. Retracted papers mentioned more frequently on Twitter/X are corrected significantly faster, while paywalled papers are corrected more slowly. This also suggests a time-sensitive design opportunity: external attention signals may be most useful as early-warning cues, and systems should surface them quickly before their association with repair fades. This pattern is consistent with the idea that social media attention may help surface retractions to editors, while a lack of open access may act as a barrier to verification and repair. Future tools might incorporate external alerting systems, integrating tweet activity and public discussions on platforms such as blogs or PubPeer regarding specific retractions into internal Wikipedia tools. This could help editors recognize emerging credibility issues even in articles that lack sustained editorial engagement.
@@ 5.2 Limitations and Future Work
Second, although our modeling identifies statistically detectable associations with correction speed, it cannot fully account for the editorial processes and sociotechnical dynamics that shape citation repair on Wikipedia. Decisions to add, retain, or remove citations are often entangled with disputes over content reliability, page notability, or conflicting interpretations of sourcing policies. Because we infer repair from revision traces, we observe outcomes more directly than the deliberative processes that produce them. Our analysis does not surface these nuanced deliberations, which may vary across topics, editor groups, or article histories. Qualitative research—such as interviews with active editors, analysis of talk page discussions, or observational studies of how citation templates are used—could provide deeper insight into how retraction cues are interpreted and acted upon in practice. More concretely, future work could examine how retraction signals travel through editors’ day-to-day workflows—especially for bot-mediated flags. For example, qualitative and trace-based mixed methods could investigate (i) how editors first notice RetractionBot templates (e.g., via watchlists, category maintenance, WikiProject backlogs, or recent-changes patrolling), (ii) how the presence of a template is translated into a specific repair action (complete removal versus in-text acknowledgment), and (iii) what points of friction or disagreement arise when deciding whether a retracted source can still be used for non-claim-supporting purposes (e.g., documenting controversy). More broadly, this line of inquiry could extend to upstream, structured sources of retraction metadata in the Wikimedia ecosystem. For instance, future work could examine how retraction status is curated and coordinated in Wikidata (e.g., through the Wikidata WikiProject on retracted research Wikidata contributors (2024)), and whether—and through what technical pathways—such structured signals propagate into Wikipedia’s citation templates, bots, and editor-facing workflows. In addition, follow-up work could trace repeated introductions of the same retracted paper across multiple articles to distinguish independent sourcing from cross-page copying or “secondary citation” practices (e.g., copying bibliographic details from other Wikipedia pages, external summaries, or locally stored PDFs). Such analyses would help clarify why a non-trivial share of citations are introduced after retraction despite publisher-level warnings, and would better ground design interventions in the realities of editors’ verification routines.
Another limitation lies in the scope of our survival modeling. While our extended Cox model with start–stop intervals and time-varying terms is effective for estimating associations and accommodating non-proportional hazards, it does not capture potentially complex, non-linear interactions among predictors. For instance, visibility signals, page structure, and retraction characteristics may combine in ways that are not well represented by linear log-hazard specifications. Future work could explore interaction-focused modeling, flexible approaches (e.g., spline-based effects or more flexible statistical models), or mixed-methods designs to uncover richer mechanisms behind citation repair.
Finally, we must acknowledge a critical limitation in our classification of “non-problematic” citations. Our definition relies on human readers’ cognitive capacity to interpret in-text warnings (e.g., “This paper was retracted”) and appropriately discount subsequent claims. However, this assumption may not hold for Large Language Models (LLMs) that utilize Wikipedia as a foundational training corpus. Recent scholarship indicates that generative AI models may fail to contextually “unlearn” retracted findings even when textual flags are present Thelwall et al. (2025). Consequently, citations that we classify as “corrected” for humans readers may still pose a persistent risk of propagating misinformation into the training data of downstream AI applications.
@@ 6 Conclusion
This study examined the presence and persistence of retracted paper citations on Wikipedia, and modeled the timing of reader-facing correction events using survival analysis. Our findings reveal that while some citations are corrected promptly, many persist uncorrected for extended periods, with a median time to correction of 3.68 years (1,344 days). In our extended Cox model, a higher hazard of correction is associated with citations flagged by RetractionBot, citations framed as controversial at the time of addition, open access availability, and greater Twitter/X attention—although the Twitter/X association is time-dependent and attenuates over time. Conversely, a high academic citation count is associated with a slower correction time, which may reflect the difficulty of challenging sources perceived as highly authoritative. We do not observe a stable average difference by retraction reason in the baseline effect; instead, the association appears time-varying, suggesting that retraction characteristics may become more predictive later in the risk period rather than uniformly across time. These findings highlight a fundamental disconnect between the public availability of retraction data and the community’s collaborative workflows, which often fail to translate this information into the reader-facing edits required for a complete repair. Addressing this challenge requires moving beyond simple error flagging toward a tighter integration of information systems and human editorial processes. Design interventions that surface different types of signals to support editorial work—such as contextual signals embedded in editing interfaces, structured data signals within citation templates, and prioritized attention signals on community dashboards—could significantly improve responsiveness to invalidated science. More broadly, our results suggest that citation repair timing covaries with visibility cues and page-level organization, pointing to opportunities to strengthen retraction legibility and routing at the point of editing while supporting sustained cross-page maintenance work. By combining better information routing with sustained human oversight, Wikipedia can more effectively manage citation-level credibility failures and strengthen its role as a reliable public knowledge platform.
@@ Appendix A Survival Analysis Model Selection Details
Table A1: First round test of the Proportional Hazards Assumption. The test is based on Schoenfeld residuals [54]. A significant p-value (p < 0.05) indicates that a variable violates the assumption.
Variable $\chi^{2}$ df p-value
RetractionBot 1.505 1 0.220
Tweet Mention Count (at risk start) 10.768 1 0.001
Open Access 13.106 1 0.0003
Team Size 0.585 1 0.444
Time to Retraction (days) 1.070 1 0.301
Pre-Retraction Editorial Event 0.103 1 0.748
Retraction Reason (salient) 16.592 1 4.6e-05
Unique Editor Count (at risk start) 4.714 1 0.030
Controversy Noted Initially 0.346 1 0.556
Media Count (at risk start) 2.118 1 0.146
Categories Count (at risk start) 10.857 1 0.001
Reference Count (at risk start) 7.441 1 0.006
Academic Citation Count (at risk start) 0.732 1 0.392
Life Sciences (domain) 33.160 1 8.5e-09
Physical Sciences (domain) 1.706 1 0.191
Social Sciences (domain) 1.090 1 0.296
GLOBAL 64.828 16 7.9e-08
Table A2: Second round test of the Proportional Hazards Assumption. The test is based on Schoenfeld residuals [54]. A significant p-value (p < 0.05) indicates that a variable violates the assumption.
Variable $\chi^{2}$ df p-value
RetractionBot 0.003 1 0.9577
Tweet Mention Count (at risk start) 10.755 1 0.0010
Open Access 5.197 1 0.0226
Team Size 0.245 1 0.6204
Time to Retraction (days) 0.299 1 0.5846
Pre-Retraction Editorial Event 0.315 1 0.5744
Retraction Reason (salient) 5.266 1 0.0218
Unique Editor Count (at risk start) 0.186 1 0.6659
Controversy Noted Initially 0.065 1 0.7981
Media Count (at risk start) 0.325 1 0.5687
Categories Count (at risk start) 2.240 1 0.1345
Reference Count (at risk start) 3.047 1 0.0809
Academic Citation Count (at risk start) 0.042 1 0.8384
GLOBAL 28.147 13 0.0086
Appendix B Independent Variable Multicollinearity
Table A3: Collinearity diagnostics based on Lee & Weissfeld (1996) for the time-dependent Cox model [71]. The analysis assesses the condition of the design matrix including the time-dependent covariate. The maximum Condition Index (CI) is 2.88, which is well below the strict threshold of 10 recommended for time-dependent models, confirming that multicollinearity is not a concern. CI diagnostics are based on model-based covariance (non-robust); robust SE are used for hypothesis tests and confidence intervals in the main model.
Variable Variance Proportion Status
(Max CI = 2.88)
Academic Citation Count (at risk start) 0.615 Safe
Media Count (at risk start) 0.235 Safe
Tweet Mention Count (at risk start) 0.103 Safe
Controversy Noted at Initial Citation 0.012 Safe
Pre-Retraction Editorial Event 0.008 Safe
Open Access 0.008 Safe
Categories Count (at risk start) 0.006 Safe
Time to Retraction (days) 0.004 Safe
Reference Count (at risk start) 0.003 Safe
RetractionBot Flagged 0.002 Safe
Retraction Reason (salient) 0.002 Safe
Number of Unique Editors (at risk start) 0.001 Safe
Team Size <0.001 Safe
Note: Multicollinearity is indicated when two or more variables have variance proportions >0.5>0.5
associated with the same large Condition Index (>10>10).
Appendix C Extended Cox Model Results
Table A4: Extended Cox proportional hazards model predicting time to correction (start–stop format). RetractionBot is modeled as a time-dependent covariate (0 pre-flag, 1 post-flag). Time-varying coefficients are modeled via covariate × log(time) interactions (denoted by “X Time”). Robust (sandwich) standard errors clustered by citation pair are reported [38].
Variable Hazard Ratio (95% CI) p-value
Paper Features
Tweet Mention Count 1.62 (1.35 – 1.94) $0.001^{***}$
     $\times$ Time 0.97 (0.95 – 1.00) $0.032^{*}$
Open Access 1.68 (1.11 – 2.55) $0.015^{*}$
     $\times$ Time 1.01 (0.95 – 1.07) 0.778
Team Size 1.13 (0.99 – 1.29) 0.061
Academic Citations (Prior Risk) 0.85 (0.73 – 0.97) $0.020^{*}$
Retraction Event Features
Time to Retraction 1.16 (0.99 – 1.37) 0.063
Pre-Retraction Editorial Event 0.98 (0.69 – 1.41) 0.921
Retraction Reason (Salient) 0.97 (0.68 – 1.40) 0.881
     $\times$ Time 1.07 (1.01 – 1.13) $0.018^{*}$
Wikipedia Page Features
RetractionBot Flagged (time-dependent covariate) 2.34 (1.91 – 2.86) $0.001^{***}$
Controversy Noted at Initial Citation 1.63 (1.03 – 2.60) $0.039^{*}$
Categories Count 1.27 (1.11 – 1.45) $0.001^{***}$
Reference Count 1.05 (0.91 – 1.20) 0.534
Unique Editors 1.05 (0.92 – 1.21) 0.436
Media Count 0.97 (0.85 – 1.10) 0.614
Significance codes: $^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$